# On variants of Chebyshev’s conjecture

### v2

@@ 1.3. Results on averages over Dirichlet characters
As naturally expected, Theorem 3 generalizes to a result for general Dirichlet $L$-functions, as stated in Theorem 8 below. This generalization is also described by inequalities for oscillating sums, however, these sums can be rewritten as those with a constant sign in both the first and second halves. For the following results, for an imprimitive Dirichlet character $\chi$, we define the GRH to hold for $L(s,\chi)$ as the GRH to hold for $L(s,\chi^{\ast})$, where $\chi^{\ast}$ is a primitive Dirichlet character inducing $\chi$.
@@
    $\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$ (1.21)
@@
    $\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}$ (1.22)
@@
    $\lim_{x\to\infty}$ $\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)\log x}\right)$ (1.23)
@@
    $\lim_{x\to\infty}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (1.24)
@@
    $\lim_{x\to\infty}$ $\left(\frac{1}{\log x}\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)(\log x)^{2}}\right)$ (1.26)
@@
    $\lim_{x\to\infty}\,\frac{1}{\log x}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$ (1.27)
@@
If $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$, then the right-hand sides of (1.26) and (1.27) vanish. Hence, the second and fourth claims of Theorem 6, as well as the third and fifth claims, are consistent.
As for (1.10), nonpositivity of (1.22) can be expressed as an inequality for nonnegative sums
@@
The previous (1.13) corresponds to the case $q=1$ of this inequality.
According to the second and third statements of Theorem 6, the sign of the sums (1.21) and (1.22) for sufficiently large $x$, which are assumed in the first statement, should be determined by the constant
@@
from (3.2), (4.11), and (4.12) below, where $C_{0}$ is the Euler–Mascheroni constant. It follows that $C(q)<0$ for all odd primes $\leq 47$, while $C(q)>0$ for all primes $\geq 53$.
The sign constancy of (1.22) is conditional, however, we find that (1.22) is unconditionally negative on average for all sufficiently large $x>1$.
@@
$\sum_{3\leq q\leq Q}$ $\left(\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (1.29)
@@
$\sum_{3\leq q\leq Q}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq xe^{2}}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}=-\frac{8}{9}\,e^{3}\,x\sqrt{x}\,(1+o(1))\quad\text{as $x\to\infty$}$ (1.30)
@@
The reason $q=1,2$ are excluded in both (1.29) and (1.30) is that, in those cases, the Dirichlet characters modulo $q$ are only the principal character. While it is certainly desirable to obtain the negativity of the left-hand side of (1.29) or (1.30) for a large constant $Q\geq 3$, we are also interested in what nontrivial conclusions can be derived when the negativity of the left-hand side of (1.29) or (1.30) is obtained under a condition weaker than $Q\geq x$ with respect to $Q$. This remains a subject for future study.
@@ 2.1. Facts on prime numbers
$\sum_{p\leq x}\frac{\log p}{p}=\log x+O(1)\quad\text{as $x\to\infty$},$ (2.1)
@@
$\psi(x)-x=\sum_{\rho}\frac{x^{\rho}}{\rho}=O(x\exp(-c\sqrt{\log x}))\quad\text{as $x\to\infty$},$ (2.2)
@@ 2.2. Riemann zeta function
$\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{n=1}^{\infty}\frac{\Lambda(n)}{n^{s}}$ (2.3)
for $\Re(s)>1$ [23, (1.1.8)]. The Riemann xi-function (1.8) is an entire function satisfying the functional equations
$\xi(s)=\xi(1-s)\quad\text{and}\quad\xi(s)=\overline{\xi(\bar{s})}$ (2.4)
@@
$\sum_{\rho}|\rho|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (2.5)
@@
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{n^{s}}$ $=\frac{x^{1-s}}{1-s}-\frac{\zeta^{\prime}}{\zeta}(s)-\sum_{\rho}\frac{x^{\rho-s}}{\rho-s}+\sum_{k=1}^{\infty}\frac{x^{-2k-s}}{2k+s}$ (2.6)
@@ 2.3. Dirichlet $L$-function
$\frac{L^{\prime}}{L}(s,\chi)=\sum_{n=1}^{\infty}\frac{\Lambda(n)\chi(n)}{n^{s}}$ (2.9)
@@
$\xi(s,\chi):=\left(\frac{q}{\pi}\right)^{(s+\kappa)/2}\Gamma\left(\frac{s+\kappa}{2}\right)L(s,\chi).$ (2.10)
@@
$\xi(s,\chi)=\varepsilon(\chi)\xi(1-s,\overline{\chi}),\quad\varepsilon(\chi):=\frac{\tau(\chi)}{i^{\kappa}\sqrt{q}}$ (2.11)
@@
$\sum_{\rho_{\chi}}|\rho_{\chi}|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (2.12)
where $\rho_{\chi}$ runs over all nontrivial zeros of $L(s,\chi)$ counted with multiplicity. By an argument similar to the proof of (2.6), we obtain
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}=-\frac{L^{\prime}}{L}(s,\chi)-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-s}}{\rho_{\chi}-s}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-s}}{2k+\kappa+s}$ (2.13)
@@
Suppose that $L(s,\chi)$ has a zero of order $m\geq 1$ at $s=1/2$. By isolating the term of $\rho_{\chi}=1/2$ in the second sum on the right-hand side of (2.13), then taking the limit $s\to 1/2$, we obtain
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ $=-m\log x-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]$ (2.14)
@@ 3.1. Proof of Theorem 1
The equality in (1.5) is obtained by integrating (1.3). First, we prove (1.5) and (1.6) assuming the RH. We set
@@
$f(x)-4\sqrt{x}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)$ (3.1)
@@
for $x\geq 2$. The second sum on the second line of the right-hand side is clearly bounded for $x\geq 2$. The first sum on the second line of the right-hand side is also bounded for $x\geq 2$, because $\rho-1/2$ is purely imaginary by the RH and (2.5). On the other hand,
@@
by the taking logarithmic derivative of the first equation in (2.4) and using $(\Gamma^{\prime}/\Gamma)(1/4)=-\pi/2-3\log 2-C_{0}=-4.22745\dots<0$ ([15, p. 411]). Therefore,
$\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)=\frac{1}{2}\left(\frac{\pi}{2}+\log 8\pi+C_{0}\right)=2.68609\dots>0.$ (3.2)
Hence, (1.5) holds for all sufficiently large $x\geq 2$. Furthermore, dividing both sides of (3.1) by $\log x$ and taking the limit as $x\to\infty$, we obtain (1.6).
By (3.2), equality (1.6) implies that (1.5) holds for all sufficiently large $x$. Therefore, it remains to show that the RH holds if there exist $x_{0}\geq 2$ such that (1.5) is valid for all $x\geq x_{0}$. The key point is the integral formula
$\int_{1}^{\infty}(-f(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1,$ (3.3)
@@
if $\Re(s)>1$, where the interchange of the order of integration and summation is justified by the absolute convergence of the series for $\Re(s)>1$ and Fubini’s theorem. Therefore, we obtain (3.3) by (2.3).
@@
By the integral formula (3.3) and $\int_{1}^{\infty}\sqrt{x}\cdot x^{-s+1/2}dx/x=1/(s-1)$ for $\Re(s)>1$,
$\int_{1}^{\infty}$ $(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x}$ (3.4)
@@
The simple pole of $(s-1/2)^{-2}(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. If we divide the left-hand side of (3.4) into
@@
the second integral defines an entire function. Applying Proposition 1 to the first integral, we find that the abscissa of convergence of this integral is at most $1/2$. Therefore, the integral converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there. This implies, by (3.4), that $(\zeta^{\prime}/\zeta)(s)$ has no poles in the right half-plane $\Re(s)>1/2$, which in turn implies that the RH holds. $\Box$
@@ 3.2. Proof of Corollary 1
Assuming the RH, it is shown that $g_{0}(t)$ is nonnegative for all sufficiently large $t>0$ by (1.9) and (3.1) as in the proof of Theorem 1. Therefore, we prove the converse. Using (1.9), (3.3), and the integral formula
@@
$\int_{0}^{\infty}(-g_{0}(t))\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\left(\frac{\zeta^{\prime}}{\zeta}(s)+\frac{1}{s-1}+\frac{1}{s}\right)$ (3.5)
for $\Re(s)>1$. The simple pole of $(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. Hence, if we use (3.5) instead of (3.4), the RH follows from the nonnegativity of $g_{0}(t)$ for all sufficiently large $t>0$ as in the proof of Theorem 1. $\Box$

### v3

@@ 1.3. Results on averages over Dirichlet characters
As naturally expected, Theorem 3 generalizes to a result for general Dirichlet $L$-functions, as stated in Theorem 8 below. This generalization is also described by inequalities for oscillating sums. By considering the averages of these sums, we obtain the following result. As in (13), this average can be rewritten as the difference of two sums, each having a constant sign in the first and second halves. For the following results, for an imprimitive Dirichlet character $\chi$, we define the GRH to hold for $L(s,\chi)$ as the GRH to hold for $L(s,\chi^{\ast})$, where $\chi^{\ast}$ is a primitive Dirichlet character inducing $\chi$.
@@
    $\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$ (21)
@@
    $\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}$ (22)
@@
    $\lim_{x\to\infty}$ $\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)\log x}\right)$ (23)
@@
    $\lim_{x\to\infty}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (24)
@@
    $\lim_{x\to\infty}$ $\left(\frac{1}{\log x}\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)(\log x)^{2}}\right)$ (26)
@@
    $\lim_{x\to\infty}\,\frac{1}{\log x}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$ (27)
@@
If $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$, then the right-hand sides of (26) and (27) vanish. Hence, the second and fourth claims of Theorem 6, as well as the third and fifth claims, are consistent.
As for (10), nonpositivity of (22) can be expressed as an inequality for nonnegative sums
@@
The previous (13) corresponds to the case $q=1$ of this inequality.
According to the second and third statements of Theorem 6, the sign of the sums (21) and (22) for sufficiently large $x$, which are assumed in the first statement, should be determined by the constant
@@
from (47), (65), and (66) below, where $C_{0}$ is the Euler–Mascheroni constant. It follows that $C(q)<0$ for all odd primes $\leq 47$, while $C(q)>0$ for all primes $\geq 53$.
The sign constancy of (21) and (22) is conditional. However, we find that they are unconditionally negative on average for all sufficiently large $x>1$.
@@
$\sum_{3\leq q\leq Q}$ $\varphi(q)\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}\right)$ (29)
@@
$\sum_{3\leq q\leq Q}\varphi(q)\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}=-\frac{8}{9}\,e^{3}\,x\sqrt{x}\,(1+o(1))\quad\text{as $x\to\infty$}$ (30)
@@
The reason $q=1,2$ are excluded in both (29) and (30) is that, in those cases, the Dirichlet characters modulo $q$ are only the principal character. While it is certainly desirable to obtain the negativity of the left-hand side of (29) or (30) for a large constant $Q\geq 3$, we are also interested in what nontrivial conclusions can be derived when the negativity of the left-hand side of (29) or (30) is obtained under a condition weaker than $Q\geq x$ with respect to $Q$. This remains a subject for future study.
@@ 2.1. Facts on prime numbers
$\sum_{p\leq x}\frac{\log p}{p}=\log x+O(1)\quad\text{as $x\to\infty$},$ (31)
@@
$\psi(x)-x=\sum_{\rho}\frac{x^{\rho}}{\rho}=O(x\exp(-c\sqrt{\log x}))\quad\text{as $x\to\infty$},$ (32)
@@ 2.2. Riemann zeta function
$\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{n=1}^{\infty}\frac{\Lambda(n)}{n^{s}}$ (33)
for $\Re(s)>1$ [23, (1.1.8)]. The Riemann xi-function (8) is an entire function satisfying the functional equations
$\xi(s)=\xi(1-s)\quad\text{and}\quad\xi(s)=\overline{\xi(\bar{s})}$ (34)
@@
$\sum_{\rho}|\rho|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (35)
@@
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{n^{s}}$ $=\frac{x^{1-s}}{1-s}-\frac{\zeta^{\prime}}{\zeta}(s)-\sum_{\rho}\frac{x^{\rho-s}}{\rho-s}+\sum_{k=1}^{\infty}\frac{x^{-2k-s}}{2k+s}$ (36)
@@ 2.3. Dirichlet $L$-function
$\frac{L^{\prime}}{L}(s,\chi)=\sum_{n=1}^{\infty}\frac{\Lambda(n)\chi(n)}{n^{s}}$ (39)
@@
$\xi(s,\chi):=\left(\frac{q}{\pi}\right)^{(s+\kappa)/2}\Gamma\left(\frac{s+\kappa}{2}\right)L(s,\chi).$ (40)
@@
$\xi(s,\chi)=\varepsilon(\chi)\xi(1-s,\overline{\chi}),\quad\varepsilon(\chi):=\frac{\tau(\chi)}{i^{\kappa}\sqrt{q}}$ (41)
@@
$\sum_{\rho_{\chi}}|\rho_{\chi}|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (42)
where $\rho_{\chi}$ runs over all nontrivial zeros of $L(s,\chi)$ counted with multiplicity. By an argument similar to the proof of (36), we obtain
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}=-\frac{L^{\prime}}{L}(s,\chi)-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-s}}{\rho_{\chi}-s}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-s}}{2k+\kappa+s}$ (43)
@@
Suppose that $L(s,\chi)$ has a zero of order $m\geq 1$ at $s=1/2$. By isolating the term of $\rho_{\chi}=1/2$ in the second sum on the right-hand side of (43), then taking the limit $s\to 1/2$, we obtain
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ $=-m\log x-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]$ (44)
@@ 3.1. Proof of Theorem 1
The equality in (5) is obtained by integrating (3). First, we prove (5) and (6) assuming the RH. We set
@@
$f(x)-4\sqrt{x}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)$ (46)
@@
for $x\geq 2$. The second sum on the second line of the right-hand side is clearly bounded for $x\geq 2$. The first sum on the second line of the right-hand side is also bounded for $x\geq 2$, because $\rho-1/2$ is purely imaginary by the RH and (35). On the other hand,
@@
by taking the logarithmic derivative of the first equation in (34) and using $(\Gamma^{\prime}/\Gamma)(1/4)=-\pi/2-3\log 2-C_{0}=-4.22745\dots<0$ ([15, p. 411]). Therefore,
$\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)=\frac{1}{2}\left(\frac{\pi}{2}+\log 8\pi+C_{0}\right)=2.68609\dots>0.$ (47)
Hence, (5) holds for all sufficiently large $x\geq 2$. Furthermore, dividing both sides of (46) by $\log x$ and taking the limit as $x\to\infty$, we obtain (6).
By (47), equality (6) implies that (5) holds for all sufficiently large $x$. Therefore, it remains to show that the RH holds if there exist $x_{0}\geq 2$ such that (5) is valid for all $x\geq x_{0}$. The key point is the integral formula
$\int_{1}^{\infty}(-f(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1,$ (48)
@@
if $\Re(s)>1$, where the interchange of the order of integration and summation is justified by the absolute convergence of the series for $\Re(s)>1$ and Fubini’s theorem. Therefore, we obtain (48) by (33).
@@
By the integral formula (48) and $\int_{1}^{\infty}\sqrt{x}\cdot x^{-s+1/2}dx/x=1/(s-1)$ for $\Re(s)>1$,
$\int_{1}^{\infty}$ $(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x}$ (49)
@@
The simple pole of $(s-1/2)^{-2}(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. If we divide the left-hand side of (49) into
@@
the second integral defines an entire function. Applying Proposition 1 to the first integral, we find that the abscissa of convergence of this integral is at most $1/2$. Therefore, the integral converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there. This implies, by (49), that $(\zeta^{\prime}/\zeta)(s)$ has no poles in the right half-plane $\Re(s)>1/2$, which in turn implies that the RH holds. $\Box$
@@ 3.2. Proof of Corollary 1
Assuming the RH, it is shown that $g_{0}(t)$ is nonnegative for all sufficiently large $t>0$ by (9) and (46) as in the proof of Theorem 1. Therefore, we prove the converse. Using (9), (48), and the integral formula
@@
$\int_{0}^{\infty}(-g_{0}(t))\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\left(\frac{\zeta^{\prime}}{\zeta}(s)+\frac{1}{s-1}+\frac{1}{s}\right)$ (50)
for $\Re(s)>1$. The simple pole of $(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. Hence, if we use (50) instead of (49), the RH follows from the nonnegativity of $g_{0}(t)$ for all sufficiently large $t>0$ as in the proof of Theorem 1. $\Box$